# Exploiting Observation Bias to Improve Matrix Completion

### v2

@@ Appendix D Imputation
This concludes the proof. ∎
Imputation.
We are now ready to present Lemma D.2 where we establish finite-sample guarantees on our matrix imputation procedure.
Lemma D.2.
Suppose that $\theta(\cdot)$, ϕ()italic-$\phi(\cdot)$, and $N_{n}$ satisfy the same assumptions as in Lemma D.1. Let $\delta\in(0,1)$. Then, the following event:
@@
$\max_{(i,j)\in\mathcal{U}^{\star}\times\mathcal{V}^{\star}}|\hat{h}_{ij}-h_{ij}|\leq c\left(\frac{d^{3}B^{2}}{(m^{2}\kappa^{4}\wedge m^{4}\kappa^{8})}\wedge 1\right)\left(LB\varepsilon_{n}+\frac{\sigma}{n}\left(\frac{2}{\varepsilon_{n}}\right)^{d-1}\sqrt{\frac{1}{\rho_{n}}\log\left(\frac{e\,n}{\delta}\right)}\right)$
@@
Step 1. (Revealed entries of interest). By Lemma D.1, the event: there exists $\mathcal{I}\subseteq\mathcal{U}^{\star}$, $\mathcal{J}\subseteq\mathcal{V}^{\star}$ such that $|\mathcal{I}|=|\mathcal{J}|=d$, and
$\forall i^{\prime}\in\mathcal{I},\forall j^{\prime}\in\mathcal{J},\qquad\bar{h}_{ij^{\prime}}\neq\star,\quad\bar{h}_{i^{\prime}j^{\prime}}\neq\star,\quad\bar{h}_{i^{\prime}j}\neq\star,\qquad\text{and}\qquad\sigma_{d}(\mathbf{H}_{\mathcal{I},\mathcal{J}})\geq c\,m^{2}\kappa^{4}$ (109)
@@
Step 2. (Error bound for revealed entries). Next, we know by Lemma C.1, that the event: for all entries $(i,j)\in\mathcal{U}^{\star}\times\mathcal{V}^{\star}$, such that $\bar{h}_{ij}\neq\star$,
$|\bar{h}_{ij}-h_{ij}|\leq\delta_{\operatorname{noise}}(n)=14LB\varepsilon_{n}+\sqrt{\frac{\sigma^{2}}{2cn^{2}\rho_{n}}\left(\frac{2}{\varepsilon_{n}}\right)^{2d-2}\log\left(\frac{2n^{2}}{\delta}\right)}$ (110)
holds with probability at least $1-\delta$, provided that $\varepsilon_{n}$ satisfies the condition (89) and $n$ is large enough as described in (90).
Step 3. (Error bound for imputed entries). Now, we are ready to use Lemma D.3, which, under the intersection of the two events described by and , entails that
$|\hat{h}_{ij}-h_{ij}|$ $\leq\left(\sqrt{2}+\frac{4\sqrt{2}d^{2}B^{2}}{\sigma_{d}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}+\frac{2(1+\sqrt{5})4d^{3}B^{2}}{\sigma_{d}^{2}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}\right)\delta_{\operatorname{noise}}(n)$ (111)
$\leq\left(\sqrt{2}+\frac{4\sqrt{2}d^{2}B^{2}}{cm^{2}\kappa^{4}}+\frac{2(1+\sqrt{5})4d^{3}B^{2}}{c^{2}m^{4}\kappa^{8}}\right)\delta_{\operatorname{noise}}(n)$ (112)

### v3

@@ Appendix B Proof of Proposition 3
This concludes the proof.
B.2 Proof of Proposition 5
B.2.1 Properties of $\theta$ and $\phi$
Lemma 5.
  • (i)
  • (ii)
Proof of Lemma 5.
To prove (i) and (ii), it suffices to establish that the mappings $\theta$ and $\tilde{\theta}$ are bi-Lipchitz with the appropriate constants, which in turn ensures the weaker statement that both mappings are homeomorphisms.
$\|\theta(\mathbf{x})-\theta(\mathbf{y})\|^{2}=e_{1}+e_{2}$ (87)
with $e_{1}=\left|\|\theta(\mathbf{x})\|-\|\theta(\mathbf{y})\|\right|^{2}$ and $e_{2}=\|\theta(\mathbf{x})\|\|\theta(\mathbf{y})\|\left\|\tilde{\theta}(\mathbf{x})-\tilde{\theta}(\mathbf{y})\right\|^{2}$. Given that $\theta(\cdot)$ is non-degenerate with parameters $m,\kappa>0$, it holds that
$\|\theta(\mathbf{x})-\theta(\mathbf{y})\|^{2}$ $\geq\|\theta(\mathbf{x})\|\|\theta(\mathbf{y})\|\left\|\tilde{\theta}(\mathbf{x})-\tilde{\theta}(\mathbf{y})\right\|^{2}$
$\geq m^{2}\kappa^{2}\|\mathbf{x}-\mathbf{y}\|^{2}.$
Thus, it must hold that $\theta$ is $(L,m\kappa)$-bilipschitz. Consequently, $\theta$ is injective and continuous and its inverse is also continuous. Thus, it is an homeomorphism.
$L^{2}\|\mathbf{x}-\mathbf{y}\|^{2}$ $\geq\|\theta(\mathbf{x})-\theta(\mathbf{y})\|^{2},$
$\|\theta(\mathbf{x})-\theta(\mathbf{y})\|^{2}$ $\geq m^{2}\|\tilde{\theta}(\mathbf{x})-\tilde{\theta}(\mathbf{y})\|^{2}$
Thus, it follows that $\tilde{\theta}$ is $(L/m,\kappa)$-bi-Lipchitz. This means that $\tilde{\theta}$ is continuous on $\mathcal{S}^{d-1}$, is bijection from $\mathcal{S}^{d-1}$ to $\tilde{\theta}(\mathcal{S}^{d-1})\subseteq\mathcal{S}^{d-1}$, and its inverse is also continuous. This means that $\tilde{\theta}$ is a homeomorphsim and it must be the case that $\tilde{\theta}(\mathcal{S}^{d-1})=\mathcal{S}^{d-1}$.
Lemma 6.
Proof of Lemma 6.
$2-2\cos(\tilde{\alpha})=\|\tilde{\theta}(\mathbf{x})-\mathbf{z}\|^{2}=\|\tilde{\theta}(\mathbf{x})-\tilde{\theta}(\mathbf{x}_{\mathbf{z}}),\|^{2}$
and
$\|\tilde{\theta}(\mathbf{x})-\tilde{\theta}(\mathbf{x}_{\mathbf{z}})\|^{2}\geq\kappa^{2}\|\mathbf{x}-\mathbf{x}_{\mathbf{z}}\|^{2}=\kappa^{2}(2-2\cos(\alpha)).$
Rearranging the terms in the above inequality we obtain
$\cos(\tilde{\alpha})\leq(1-\kappa^{2})+\kappa^{2}\cos(\alpha),$
which leads to $\tilde{\alpha}\geq\arccos(1-\kappa^{2}+\kappa^{2}\cos(\alpha))$ (because $\cos(\cdot)$ is a strictly decreasing function from $[0,\pi]$ to $[-1,1]$, and $\alpha,\tilde{\alpha}\in[0,\pi]$).
B.2.2 Properties of high-dimensional simplices
The following result on regular simplicies is borrowed from [31] (see their Theorem 1).
Lemma 7.
$\langle\mathbf{a}_{i},\mathbf{a}_{j}\rangle=\frac{-1}{d-1}.$
Lemma 8.
$\sigma_{1}(\mathbf{X})$ $=\sqrt{1+(d-1)\left(1-\frac{c^{2}}{2}\right)},$
$\quad\sigma_{2}(\mathbf{X})$ $=\dots=\sigma_{d}(\bm{X})=\frac{c}{\sqrt{2}}.$
Proof of Lemma 8.
Let us start by observing that, for all $i,j\in[d]$,
$(\mathbf{X}^{\top}\mathbf{X})_{ij}=\begin{cases}1&\text{ if }i=j,\\ (2-c^{2})/2&\text{ if }i\neq j.\end{cases}$
$\mathbf{z}_{1}^{\top}$ $=\begin{bmatrix}1&1&1&\cdots&1\end{bmatrix},$
$\mathbf{z}_{2}^{\top}$ $=\begin{bmatrix}-1&d-1&-1&\cdots&-1\end{bmatrix},$
$\mathbf{z}_{3}^{\top}$ $=\begin{bmatrix}-1&-1&d-1&\cdots&-1\end{bmatrix},$
$\vdots$
$\mathbf{z}_{d}^{\top}$ $=\begin{bmatrix}-1&-1&-1&\cdots&d-1\end{bmatrix},$
and for all $i\in\{2,\dots,d\}$, we have
$(\mathbf{X}^{\top}\mathbf{X})\mathbf{z}_{i}=(c^{2}/2)\mathbf{z}_{i}.$
This proves that $\mathbf{X}^{\top}\mathbf{X}$ has two eigenvalues such that the smallest one $(c^{2}/2)$ has multiplicity $d-1$. We conclude that $\sigma_{1}(\mathbf{X})=\sqrt{(1+(d-1)(2-c^{2})/2)},\quad\text{and}\quad\sigma_{2}(\mathbf{X})=\cdots=\sigma_{d}(\mathbf{X})=c/\sqrt{2}$
B.2.3 Proof of Proposition 5
Algorithm 1 Search rank $d$-sub-matrix in $\mathbf{X}$
Input: $\mathbf{u},\mathbf{v}\in\mathcal{S}^{d-1}$
Step 1: Finding sufficiently close features
Find the geodesic $\Gamma$ between $\mathbf{u}$ and $\mathbf{v}$ on the hypersphere $\mathcal{S}^{d-1}$
Find the midpoint $\mathbf{m}\in\mathcal{S}^{d-1}$ of $\Gamma$
Find the point $\mathbf{x}$ on $\Gamma$ such that $\arccos(\langle\mathbf{x},\mathbf{m}\rangle)=\pi/8$ and is closer to $\mathbf{v}$
Find the point $\mathbf{y}$ on $\Gamma$ such that $\arccos(\langle\mathbf{y},\mathbf{m}\rangle)=\pi/8$ and is closer to $\mathbf{u}$
Let $\mathcal{C}(\mathbf{x},\pi/12)$ be the spherical cap on $\mathcal{S}^{d-1}$ specified by the point $\mathbf{x}$ and polar angle $\pi/12$
Let $\mathcal{C}(\mathbf{y},\pi/12)$ be the spherical cap on $\mathcal{S}^{d-1}$ specified by the point $\mathbf{y}$ and polar angle $\pi/12$
Step 2: Finding well separated features
Let $\alpha=\arccos(1-\kappa^{2}+\kappa^{2}\cos(\pi/12))$
Let $\mathcal{C}(\tilde{\theta}(\mathbf{x}),\alpha)$ be the spherical cap on $\mathcal{S}^{d-1}$ specified by the point $\tilde{\theta}(\mathbf{x})$ and polar angle $\alpha$
Let $\mathcal{C}(\tilde{\phi}(\mathbf{y}),\alpha)$ be the spherical cap on $\mathcal{S}^{d-1}$ specified by the point $\tilde{\phi}(\mathbf{y})$ and polar angle $\alpha$
Let $\mathbf{z}_{1},\dots,\mathbf{z}_{d}$ be vertices of a regular $(d-1)$-simplex inscribed in $\partial\mathcal{C}(\tilde{\theta}(\mathbf{x}),\alpha)$
Let $\mathbf{w}_{1},\dots,\mathbf{w}_{d}$ be vertices of a regular $(d-1)$-simplex inscribed in $\partial\mathcal{C}(\tilde{\phi}(\mathbf{y}),\alpha)$
For all $i\in[d]$, set $\mathbf{x}_{i}=\tilde{\theta}^{-1}(\mathbf{z}_{i})$ and $\mathbf{y}_{i}=\tilde{\phi}^{-1}(\mathbf{w}_{i})$
Proof of Claim 3. First, we note that Lemma 5 and Lemma 6 ensure that $\tilde{\theta}$ and $\tilde{\phi}$ are well defined, and that $\mathcal{C}(\tilde{\theta}(\mathbf{x}),\alpha)\subseteq\tilde{\theta}(\mathcal{C}(\mathbf{x},\pi/12))$ and $\mathcal{C}(\tilde{\phi}(\mathbf{y}),\alpha)\subseteq\tilde{\phi}(\mathcal{C}(\mathbf{y},\pi/12))$. This in turn implies that $\mathbf{x}_{1},\dots,\mathbf{x}_{d}\in\mathcal{C}(\mathbf{x},\pi/12)$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{d}\in\mathcal{C}(\mathbf{y},\pi/12)$. It remains to lower bound the the smallest singular value of $\begin{bmatrix}\theta(\mathbf{x}_{1})&\ldots&\theta(\mathbf{x}_{d})\end{bmatrix}^{\top}\begin{bmatrix}\phi(\mathbf{y}_{1})&\ldots&\phi(\mathbf{y}_{d})\end{bmatrix}$. Under assumption A2, we have
$\qquad\qquad\qquad\qquad\geq m^{2}\inf_{\mathbf{z}\in\mathcal{S}^{d-1}}\sum_{i=1}^{d}|\tilde{\theta}(\mathbf{x}_{i})^{\top}\mathbf{z}|^{2}$
$\qquad\qquad\qquad\qquad\geq m^{2}\sigma_{d}\left(\begin{bmatrix}\tilde{\theta}(\mathbf{x}_{1})&\ldots&\tilde{\theta}(\mathbf{x}_{d})\end{bmatrix}\right)^{2}$
Therefore, we may apply Lemma 8 and obtain:
$\sigma_{d}\left(\begin{bmatrix}\theta(\mathbf{x}_{1})&\ldots&\theta(\mathbf{x}_{d})\end{bmatrix}^{\top}\begin{bmatrix}\phi(\mathbf{y}_{1})&\ldots&\phi(\mathbf{y}_{d})\end{bmatrix}\right)$
$\qquad\qquad\geq\frac{m^{2}(1-\cos^{2}(\alpha))}{\sqrt{2}}$
$\qquad\qquad=\frac{m^{2}\left(1-\left(1-\kappa^{2}+\kappa^{2}\cos(\pi/12)\right)^{2}\right)}{\sqrt{2}}$
$\qquad\qquad\geq\frac{m^{2}\kappa^{4}(1-\cos(\pi/12))\cos(\pi/12)}{\sqrt{2}}.$
Indeed, for all $i\neq j$, we have by construction that $\|\tilde{\theta}(\mathbf{x}_{i})-\tilde{\theta}(\mathbf{x}_{j})\|_{2}=\sqrt{1-\cos^{2}(\alpha)}$. Then, using the fact that $\widetilde{\theta}(\cdot)$ is $L$-lipschitz which comes from Lemma 5, we can immediately lower bound $\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}$ as claimed. The claim for $\mathbf{y}_{1},\dots,\mathbf{y}_{d}$ follows similarly.
B.2.4 Smallest singular value under perturbation
The Proposition 5 guarantees existence of $\mathbf{x}_{1},\dots,\mathbf{x}_{d}$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{d}$ in $\mathcal{S}^{d-1}$ such that $\sigma_{d}\left(\begin{bmatrix}\theta(\mathbf{x}_{1})&\ldots&\theta(\mathbf{x}_{d})\end{bmatrix}\right)\geq C$ and $\sigma_{d}\left(\begin{bmatrix}\phi(\mathbf{y}_{1})&\dots&\phi(\mathbf{y}_{d})\end{bmatrix}\right)\geq C$ for some universal positive constant $C>0$. However, we may only have approximate values of the vectors $\mathbf{x}_{1},\dots,\mathbf{x}_{d}$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{d}$. Below, we present a result showing that this does not matter too much.
Lemma 9.
Let $\mathbf{x}_{1},\dots,\mathbf{x}_{d}$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{d}$ in $\mathcal{S}^{d-1}$ such that $\sigma_{d}\left(\theta(\mathbf{X})^{\top}\phi(\mathbf{Y})\right)\geq C.$ Let $\mathbf{u}_{1},\dots,\mathbf{u}_{d}$ and $\mathbf{v}_{1},\dots,\mathbf{v}_{d}\in\mathcal{S}^{d-1}$ be such that $\|\mathbf{U}-\mathbf{X}\|_{2\to\infty}\leq\varepsilon$ and $\|\mathbf{V}-\mathbf{Y}\|_{2\to\infty}\leq\varepsilon$ for some $\varepsilon<C/(4BL)$. Then, we have
Proof of Lemma 9
As $\theta$ and $\phi$ are $L$-Lipschitz, we know
Thus, applying Weyl’s inequality, we have
$\left\lvert\sigma_{d}(\theta(\mathbf{U})^{\top}\phi(\mathbf{V}))-\sigma_{d}(\theta(\bm{X})^{\top}\phi(\bm{Y}))\right\rvert$
$\qquad\qquad\leq\sigma_{1}(\sigma_{d}(\theta(\mathbf{U})^{\top}\phi(\mathbf{V}))-\theta(\mathbf{X})^{\top}\phi(\mathbf{Y}))$
$\qquad\qquad\leq\|\theta(\mathbf{U}_{n})^{\top}\phi(\mathbf{V}_{n})-\theta(\mathbf{X})^{\top}\phi(\mathbf{Y})\|_{\max}$
$\qquad\qquad\leq\max_{i,j\in[d]}\left\lvert\theta(\mathbf{u}_{i})^{\top}\phi(\mathbf{v}_{j})-\theta(\mathbf{x}_{i})^{\top}\phi(\mathbf{y}_{j})\right\rvert$
$\qquad\qquad\leq 2BL\varepsilon$
If $\varepsilon\leq C/(4BL)$, then we have
Appendix C Proof of Proposition 4
We are now ready to prove Proposition 4. Under the assumptions made in the proposition, we establish that the following event:
@@ Appendix C Proof of Proposition 4
$\max_{(i,j)\in\mathcal{U}^{\star}\times\mathcal{V}^{\star}}|\hat{h}_{ij}-h_{ij}|\leq c\left(\frac{d^{3}B^{2}}{(m^{2}\kappa^{4}\wedge m^{4}\kappa^{8})}\wedge 1\right)$
×$\qquad\qquad\times\left(LB\varepsilon_{n}+\frac{\sigma}{n}\left(\frac{2}{\varepsilon_{n}}\right)^{d-1}\sqrt{\frac{1}{\rho_{n}}\log\left(\frac{e\,n}{\delta}\right)}\right)$
@@
Step 1. (Revealed entries of interest). By Proposition 3, the event: there exists $\mathcal{I}\subseteq\mathcal{U}^{\star}$, $\mathcal{J}\subseteq\mathcal{V}^{\star}$ such that $|\mathcal{I}|=|\mathcal{J}|=d$, and $\sigma_{d}(\mathbf{H}_{\mathcal{I},\mathcal{J}})\geq c\,m^{2}\kappa^{4}$
$\forall i^{\prime}\in\mathcal{I},\forall j^{\prime}\in\mathcal{J},\ \ \bar{h}_{ij^{\prime}}\neq\star,\ \ \bar{h}_{i^{\prime}j^{\prime}}\neq\star,\ \ \bar{h}_{i^{\prime}j}\neq\star,$ (92)
@@
Step 2. (Error bound for revealed entries). Next, we know by Proposition 2, that the event: for all entries $(i,j)\in\mathcal{U}^{\star}\times\mathcal{V}^{\star}$, such that $\bar{h}_{ij}\neq\star$,
$|\bar{h}_{ij}-h_{ij}|\leq\delta_{\operatorname{noise}}(n)$
$\qquad\triangleq 14LB\varepsilon_{n}+\sqrt{\frac{\sigma^{2}}{2cn^{2}\rho_{n}}\left(\frac{2}{\varepsilon_{n}}\right)^{2d-2}\!\!\!\log\left(\frac{2n^{2}}{\delta}\right)}$ (93)
holds with probability at least $1-\delta$, provided that $\varepsilon_{n}$ satisfies condition (16) and $n$ is such that $n\gtrsim d\log\left(e\,n/\delta\right)$.
Step 3. (Error bound for imputed entries). Now, we are ready to use Lemma 19, which, under the intersection of the two events described by and , entails that
$|\hat{h}_{ij}-h_{ij}|$
$\ \leq\left(\sqrt{2}+\frac{4\sqrt{2}d^{2}B^{2}}{\sigma_{d}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}+\frac{2(1+\sqrt{5})4d^{3}B^{2}}{\sigma_{d}^{2}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}\right)\delta_{\operatorname{noise}}(n)$
$\ \leq\left(\sqrt{2}+\frac{4\sqrt{2}d^{2}B^{2}}{cm^{2}\kappa^{4}}+\frac{2(1+\sqrt{5})4d^{3}B^{2}}{c^{2}m^{4}\kappa^{8}}\right)\delta_{\operatorname{noise}}(n)$